OpenStudy (anonymous):

simplify the expression 5e^-2x/e4x times (e^-2x)^2/10e^-x

OpenStudy (dumbcow):

whenever the base is the same and you are multiplying or dividing you can add or subtract the exponents 5e^-2x/e^4x = 5e^(-2x-4x) = 5e^-6x (e^-2x)^2 = e^-2x*e^-2x = e^(-2x-2x) = e^-4x e^-4x/10e^-x = e^-4x--x/10 = e^-3x/10 5e^-6x*e^-3x/10 = 5/10*e^(-6x-3x) = e^-9x/2

OpenStudy (dumbcow):

it could also be written as 1/(2e^9x)

OpenStudy (anonymous):

thankyou

OpenStudy (dumbcow):

your welcome